A water heater is operated by direct solar power. If the solar collector has an area of 12m2, and if the intensity delivered by sunlight is 620W/m2, how long does it take to increase the temperature of 1.8m3 of water from 21°C to 46°C? Take the water specific heat capacity to be 4186 J/(oC kg). Take the density of water to be 1g/cm. State your answer in minutes to the nearest minute. Do not enter the units.


1
Expert's answer
2021-10-05T10:08:13-0400

"AP=Q/t\to Q=APt"


"Q=cm(t_2-t_1)=c\rho V(t_2-t_1)"


"APt=c\rho V(t_2-t_1)\to t=c\rho V(t_2-t_1)/(AP)=4186\cdot1000\cdot1.8\cdot(46-21)/(12\cdot620)="


"=2532\ (s)=42.2\ (min)" . Answer